OpenStudy (anonymous):

find the focus of the parabola (x-5)^2=8(y+4)

OpenStudy (jdoe0001):

well, keeping in mind that the "focus form of a parabola" is => \(\bf (x-h)^2=4p(y-k); \pm p\) p = distance of the focus from the (h, k) vertex now look at your equation => \(\bf x-5)^2=8(y+4)\) you're pretty much given the (h, k) of the parabola it has an squared "x" component, so it's opening vertically the "4P" term is positive, so that means is opening upwards over the y-axis

OpenStudy (jdoe0001):

if you want the "p" distance, well 4P = 8 solve for "p"

OpenStudy (jdoe0001):